jigglypuff314 (jigglypuff314):

A particle moves alone the x-axis so that its acceleration at any time t≥0 is given by a(t)=2t-6. The initial velocity of the particle is -27 and at time t=1, its position is 20. For what values of t, 0≤t≤3, is the particle's instantaneous velocity the same as its average velocity on the closed interval [0,3]?

jigglypuff314 (jigglypuff314):

I has got s(t) = (1/3)t^3 - 3t^2 - 27t + 149/3 and v(t) = t^2 - 6t - 27 which meh teacher said was right, but idk what to do with themz now XD

jigglypuff314 (jigglypuff314):

oh and no calculator question... so gimme a while XD

OpenStudy (tkhunny):

Let's see... \(a(t) = 2t - 6\) \(v(t) = t^{2} - 6t + v_{0}\) We are given \(v_{0} = -27\) \(v(t) = t^{2} - 6t - 27\) \(s(t) = (1/3)t^{3} - 3t^{2} - 27t + s_{0}\) We are given: \(s(1) = (1/3) - 3 - 27 + s_{0} = 20 \implies s_{0} = (149/3)\) \(s(t) = (1/3)t^{3} - 3t^{2} - 27t + 149/3\) Very good. Now the question... For what values of t, 0≤t≤3, is the particle's instantaneous velocity the same as its average velocity on the closed interval [0,3]? How do we find the Average Velocity on [0,3]? \(\dfrac{v(3) - v(0)}{3-0}\)? What say you? Is that the average velocity or is it something else?

jigglypuff314 (jigglypuff314):

I did v(0) and v(3) tkhunny but got it wrong. I'm guessing that was avg accel

jigglypuff314 (jigglypuff314):

so I has to do s(0) and s(3) ... s(3)... 0.o no calculator...

jigglypuff314 (jigglypuff314):

@nincompoop do I get that avg velocity = 1/3 ? :3

OpenStudy (tkhunny):

Why did you do that? You didn't answer the question. Don't do it unless you KNOW what it is. Average velocity is \(\dfrac{How\;far\;did\;we\;get?}{How\;long\;did\;it\;take?} = \dfrac{s(3)-s(0)}{3-0} = \dfrac{}{}\) Okay, what do you get for Average Velocity?

jigglypuff314 (jigglypuff314):

1/3 ?

OpenStudy (tkhunny):

One piece at a time. s(3) = ?? s(0) = ?? Please provide these values. No guessing.

jigglypuff314 (jigglypuff314):

152/3 and 149/3 so avg velocity = 1/3

OpenStudy (tkhunny):

Okay, s(3) isn't working. Please give that another go. Be more careful, this time.

jigglypuff314 (jigglypuff314):

(1/3)(3^3) - 3(3^2) - 27(3) + 149/3 (27/3) - 27 - 81 + 149/3 9 - 108 + 149/3 - 99 + 149/3 -297/3 + 149/3 -148/3 ? so I did 99 times 3 = 307 wrong? XD no calculator sucks, I can't math >,<

OpenStudy (tkhunny):

That's better. So the average velocity is? (Don't forget to divide by (3-0)).

jigglypuff314 (jigglypuff314):

148+149 = 297 297/3 = 99 so -99?

OpenStudy (tkhunny):

You didn't read inside those last parentheses. \(\dfrac{s(3) - s(0)}{3 - 0}\). There is a Denominator, there.

jigglypuff314 (jigglypuff314):

oh, so 99/3 = 33 so -33?

OpenStudy (tkhunny):

Ding!!! Okay, off to part II. Find every time t in [0,3] where the Velocity actually takes on this value.

jigglypuff314 (jigglypuff314):

so then I look for -33 = (t-9)(t+3) ?

OpenStudy (tkhunny):

I suppose, but I have no idea why you bothered to factor it. Factoring does little good unless the other side of the equation is ZERO!

jigglypuff314 (jigglypuff314):

oh yeah... so -33 = t^2 - 6t - 27 then t^2 - 6t + 6 = 0 ?

OpenStudy (tkhunny):

Okay, two more things to do.

jigglypuff314 (jigglypuff314):

no no no not quadratic equation, please not quadratic equation i don't wanna D': I had it programed into meh calculator and everything so I wouldn't have to memorize it >,<

jigglypuff314 (jigglypuff314):

I'm hopeless aren't I >,<

OpenStudy (tkhunny):

Well, Complete the Square, then.

jigglypuff314 (jigglypuff314):

how would completeing the square help me? (t-3)^2 -3 = 0

OpenStudy (tkhunny):

Where do you think the Quadratic Formula came from? Solve for t.

jigglypuff314 (jigglypuff314):

oh... (t-3)^2 = 3 then t-3 = √3 & -√3 t = 3 + √3 and 3 - √3 ?

OpenStudy (tkhunny):

Okay, now answer the question. What t on [0,3] have we just found?

jigglypuff314 (jigglypuff314):

oh 3 - √3

OpenStudy (tkhunny):

Done!

jigglypuff314 (jigglypuff314):

OMG THANK YOU!!! :D

OpenStudy (tkhunny):

Free Advice: Get that quadratic formula stuck in your head. Completing the square is fine, and that IS where the formula came from, but a little better speed and accuracy can be achieved with the formula well-memorized.

OpenStudy (tkhunny):

BTW Psyduck rules!!!

OpenStudy (nincompoop):

[-B +/- SQRT(B^2 -4AC)]2A

jigglypuff314 (jigglypuff314):

ahahahaha XD thanks tkhunny!

OpenStudy (nincompoop):

if you can remember discriminant you're halfway done